Please help,
Evaluate each function for the value specified.
f(x) = x^2 − 7x + 10; Find (a) f(0), (b) f(5), and (c) f(-2).
Thanks
JMS

Answer by malakumar_kos@yahoo.com(315)


Evaluate each function for the value specified.
f(x) = x^2 − 7x + 10; Find (a) f(0), (b) f(5), and (c) f(-2).
f(x) = x^2-7x+10
a) f(0) = 0^2-7(0)+10 = 0-0+10 = 10
b) f(5) = 5^2-7(5)+10 = 25-35+10 = 35-35 = 0
c) f(-2) = (-2)^2-7(-2)+10 = 4+14+10 = 28
hence values are f(0) = 10, f(5) = 0 & f(-2) = 28